Status Report

ISS MER Status Report 9 Mar 2001

By SpaceRef Editor
March 9, 2001
Filed under ,

Excerpt from 9 March 2001: Felicity Vol. 50, ISS Program Office (Acrobat 2.1MB)

CCS HRDL MSD

At 048/ 09: 10: 58 GMT, C& C1 MDM indicated a CCS HRDL MSD Access to ggle from true to
FALSE. The crew attempted to reset per onboard procedures but failed. The ground
issued a HRDL reset and successfully cleared the indication. Crew indicated the PCS
was back to normal. CCS and PCS personnel are performing testing to attempt to
duplicate the observed conditions. Also, ODIN has asked the crew to dump the PCS
log file for analysis. Condition had been seen at KSC, but was thought to be fixed.

Lab C& C MDM- 2 Failure Indication.

At 055: 02: 33: 33 the backup C& C MDM went offline. About two minutes later it came back
online. The Bit Summary table indicates a watchdog timer error. The CCS software
does not allow the MDM to come back up on it’s own. CCS software does not provide
for this to occur since retry and recovery are inhibited. Various logs have been
downlinked and are currently being analyzed.

INT 2 MDM Bus Exception Error.

At 055: 13: 43: 46 the INT SYS MDM Bit summary table A indicated a SX Bus Exception
error. At the time the following events were occurring:

1) The station crew were entering the station after the Soyuz redocking to the FGB Nadir
Port.

2) N1RS4B- C Health status fla g set.

This condition was seen during MEIT at the Cape. IPR 65 and PRACA 2325 were opened
on this. PR# 19532 document that INT- 1 showed this failure indication.

C& C- 2 MDM Switched Over to Backup

At 22: 31: 27 (GMT) CC- 2 was running as Primary and failed to diagnostics. CC- 1 came up
as Primary. We dumped additional data to determine the exact type of ADA exception.
C& C2 has been recovered to backup.

CCS- 2 MSD Year incorrect for Filename: LDF. txt

When using the SOSMARTT application to verify files on the on- board MSDs it was
noticed that CC- 2 MSD shows a date of 1954 for the LDF. txt. We know this file is the
correct version because it was transferred from CC- 1 MSD during 5A. We would like
to investigate why the year shows incorrectly.

PCU Tubes Never Reached Temp When Heaters Were Active.

For both PCU 1 & 2, the Tube Temperatures are several degrees below their expected value.
The tube temperatures should have reached operating temperature and started cycling
between 32 and 42 deg C, the heater ON / OFF setpoints. Since the tubes have never reached
the Heater Off temperature, they have remained ON constantly, providing 15W of heat to the
tubes. The temperature must be above 20.5 for the PCU to operate normally. If the tube
temperature drops below 20.5 de g C, and the PCU is in discharge mode, the Low Temperature
FDIR will activate and turn the PCU off. To prevent this from occuring , the Low Temperature
FDIR has been Inhibited since EVA #2 on Fli ght 3A. Currently any time the vehicle is is
negative betas, it is expected that both PCUs will need to be in discharge mode to maintain
operational/ standby temperatures on the tubes (> 16.6C). The Thermal Model was skewed
“HOT” because of the implications of overheating/ overpressurizing the Tank. Revised
analysis is showing that either there is an unexpected heat sink (pulling heat away from the
tubes) or the tube heaters are undersized. Short Term: Maintain both PCUs in Discharge
Mode as required to maintain tube temperature above 16.6C (referred to as Option #1). This
generally means any time the ISS is in negative beta angles.

Four Long Term Solutions are available, and are being pursued by the ART. These options
include #2) Use Cathode heater to warm PCU tubes in case of severe need (i. e. First PCU
failure or Emer gency EVA) #2.1) Use Cathode Heater to warm PCU Tubes when Ignition
necessary. 3) Certify the PCU for liquid Xe and 4) Increase tube heater size.


FPP Failure To Acquire Data.

The FPP transferred data nominally from the Probe to the Network Control Unit (NCU) during
Flight 4A mated operations. Upon the Orbiter departure, several attempts at transferring FPP
data went astray due to the UOP 1 Failure and UOP 2 GFCI Trips. After developing a successful
plug- in- plan removing the GFCI issue, transferring the FPP data was attempted again.
Unfortunately, no communication could be established between the FPP NCU and the FPP
Probe. Loss of FPP would result is loss of ISS voltage potential data only. The IMC/ Program
Office folks responsible for the FPP developed a new version of FPP software to reduce several
networking requirements/ protocols. This software is the application that runs on the SSC
computer that is connected to the FPP NCU via parallel cable.

LAB LTA Heater not cycling as expected.

During LTA phase, the LTA heaters have not cycled off as expected. Both APCUs
appear to be supplying power to the LAB LTA heaters. It was later confirmed that
both strings of Lab LTA heaters were on from MET 3.5 hrs (time of APCU activation)
until MET 64.2 hrs. An on- orbit test of power- cycling APCU- 1 was done about 1 hour
before LAB unberth from Shuttle. The heaters had been on prior to the test. After the
power- cycle (on/ off/ on), all heaters remained OFF. After connection to Node 1, the
same three thermostats which had been powered by APCU- 1, as well as a fourth (that
had previously been unused on- orbit), had Node 1 power applied. A gain none of the
4 came ON. A re- applications of Node 1 power done on March 5, 2001 resulted in 2
thermostats turning ON, 1 staying OFF, and the fourth one tripping its upstream
RPCM output. The first impact would be the PPRV popping. Structures and
Mechanisms has reviewed the Lab pressure shell structural capability at the
increased temperature. The slight reduction in material strength capability is covered
by analysis safety factors and there are no structural concerns. Post- flight
predictions by STAR indicate that no Lab components violated upper temperature
limits during the launch- to- activation phase, as documented in LTA Heater ART
presentation “Thermal Model/ Predictions Investigation” dated 2/ 21/ 01, pgs. 4- 5. The
same type of thermostat is used to control the Lab condensation heaters. Monitoring
of the Lab condensation heater performance indicates that these heaters are cycling
on/off as expected. Cause of the problem is still unknown.

RPCM N1 4B C Health Status Flag Set

  • @ GMT 055: 13: 42: 36 N14B- C RPCM Health Flag set. Bitsumm and Fwstat BDTs were
    pulled which indicated a flipped bit in the Primary Instruction Memory in block 11.
    Block 11 was dumped, Block 11 of RPCM N1- 4B- C contains the following functions:
  • 1) Trip functions (Current and undervoltage trips)
  • Undervolta ge Recovery – May close undesired RPC channels.
  • Undervoltage trip – may cause undesirable RPC shutdown (low probability)
  • Current – May cause nuisance tripping for false over- current condition.
  • For real over- current condition, hardware function will still trip open RPCs.
  • 2) Command Processing
  • Handles receive commands (from ground) and transmit commands (from telemetry).
  • Could misinterpret one command for another. Investigation continues.

    RPC tripped during LTA Heater Test

    The LTA heater known as 4A on the lab (N1- RS1- B, sw 15) tripped during the test to
    determine heater functionality. No other test is planned right now.

    CMG 2 Loss of Comm.

    CMG 2 was “safed” by the flight software after loss of comm between the FWC and
    the GN& C MDM was detected. Bus switching FDIR and FWC power cycle were
    performed automatically. The system continued to control with 3 CMGs while CMG 2
    was recovered. Four CMGs are currently in use. Investigation continues.

    CMG spin bearing heater fault and EA fault Isolation

    CMGs are expected to show occasional setting of the EA Fault Isolation Status Bit,
    usually due to the setting of the Spin Bearing Heater fault bit. What we are seeing is
    CMG2 periodically exhibiting a Spin Motor current spike associated with these events
    and the setting of the SM PS Trip flag. Causes are being investigated but as of 02/ 28,
    no definitive cause can be identified nor can a definitive correlation with CMG
    operations or environment be identified. We will continue to monitor and assess.

    OCA Headset Cable Failure

    Crew OCA headset can listen but not talk. Problem isolated to OCS headset cable.
    Troubleshooting of cable indicated no connectivity in microphone leg of cable,
    isolated to microphone jack itself. Also, per Flight Note OPFN961, the Crew has also
    reported problems in earphone leg. The Crew performed tests and attempted repairs
    as noted in Flight Note OPFN961. Apparently the mic side was repaired to
    functionality by splicing a jack from a Sony CD speaker. Then it was discovered that
    the Earphone cable was in operative. Not having the parts to do a repair to the
    Earphone side, the Crew used the laptop external speakers. This gives the Crew a
    patched system that is presently providing voice transmit and receive. A new cable
    will be flown on flight 5A. 1. FIAR is being generated. L& M is to find out if cable was
    returned to ground.

    Periodic Annunciation of S- Band XPDR Non- Critical Failure Advisories

    The Cato team was informed by the ODIN team of periodic annunciations of Event
    Code 5106 S- Band XPDR Non- Critical Failure advisories. Data provided by ODIN
    shows the advisories being generated at 5 to 20 minute intervals with consistent in-
    alarm periods of 10 seconds. S- Band XPDR Non- Critical Failure advisory was
    inhibited at GMT 046/ 19: 42: 16. C& T has a Chit ISS0151 to work this issue.

    S-Band BSP Loss of Forward Link Audio Channel RELP Sync

  • At GMT 2001/ 050: 04: 06: 47, Audio Channel 1 RELP sync was lost. Audio Channel 1
    RELP sync returned at 050/ 04: 07: 07 and became intermittent until it was permanently
    lost at 050/ 04: 09: 57. Audio Channel 2 RELP sync was lost at 050/ 04: 07: 47.

    IMPACTS:

  • Loss of Audio Channel 1 and 2 RELP sync resulted in loss of forward link voice
    capability via both ISS S- Band Space to ground audio channels from 050/ 04: 09: 57 to
    050/ 06: 42: 37.
  • CATO performed the following actions in attempting to recover uplink voice
    capability:
  • 1. Disabled and re- enabled uplink voice outputs MCC – No Joy
  • 2. Allowed S- Band system to perform forward link acquisition sequence on nest
    TDRS – No Joy
  • 3. Commanded S- Band string to HDR Configuration – No joy
  • 4. Commanded Audio CH1 to OFF, then back to ON – Audio Ch1 RELP sync recovered
    (GMT 050/ 06: 38: 36)
  • 5. Commanded Audio Ch2 to OFF, then back to ON – Audio CH2 RELP sync recovered
    (GMT 050/ 06: 42: 37)
  • 6. Successfully performed S/ G voice check with ISS Crew on S/ G 1

    Echo on S Band (Ground).

  • “When CAPCOM is talking to the crew in the SM, the ground hears an echo. The
    ground hears a repeat of their voice. The ISS audio system was configured with S/ G 1,
    ATU Lab 1, ATU Lab2, and RSA 1 in Public Loop1 and . It is probable that the SM
    audio system was configured in a way that the voice was looped back down the S/ G
    channel.
  • Forward plan for closure:
  • CATO put to gether OCA_ 1578 US/ RS Comm Configuration procedure to configure the
    SM comm panels to prevent the audio from being looped into the S/ G channel.
  • S- BAND Gimbal Temperatures.
  • “Fli ght sensor data for the SASA RFG EL and AZ gimbals show temperatures that are
    64 De g F (35 DegC) to 104 DegF (58 DegC) different from predictions. This behavior
    has been observed since during Flight 5A to date. However sensor temperature data
    is well within maximum temperature requirement of 169 De gF (76 DegC)
  • Predicted temperatures ran ged between -2 Deg F (- 16.6 DegC) to 16 DegF (- 8.8 DegC)
    for both gimbals. Flight data temperatures range between 81.5 DegF (27.5 DegC) to
    118 De gF (48 DegC) for EL gimbal and 98 DegF (36.5 DegC) to 108 DegF (42 DegC) for
    the AZ gimbal”

    EVA Tool does not fit PCA Nonpropulsive Vent bolt head

    EVA tool does not fit ACS NPV bolt head. Investigation is underway to determine how
    this was not detected on ground. KSC engineering was requested to take the spare
    vent cover at KSC and perform a precision measurement of the bolt heads. The
    actual size was measured to be 3/ 8 inch. MOD EVA was notified of the size, the crew
    was advised to obtain the correct socket from the Node Bag on orbit, and the task
    was reattempted and successful. Fit checks completed for future flights and have
    been reviewed for redlines to the procedures. Procedural training will ensure
    corrections made during a fit check are verified against the current drawings.

    SASA Connector P4 O- Ring Dislodge

    During the EVA install of the SASA an O- Ring became dislodged from the P- 4
    connector. Request that an inspection/ evaluation of the recovered O- Ring by the
    vendor of the connector to determine the cause and possible corrective action to
    prevent future occurrences.

    Camera Bracket Sticking in Swing Arm Fitting

    “EV2 had difficulty removing the camera bracket bayonet from the MWS swing arm
    fitting.

    Treadmill Slats Broken

    On 2- 25- 01, the crew performed OCA_ 1623 TVIS Belt Slat Replacement Procedure but
    did not replace the foot plate (aft top plate). It was determined that, for stability and
    safety reasons, this foot plate needs to be re- installed. The crew is scheduled to
    complete a second TVIS IFM (part 2) (OCA_ 1759) this week to install the foot plate and
    ti ghten the two aft stabilizer bolts. This is not an immediate safety hazard so the crew
    can continue to exercise on the TVIS.

    CREW EQUIP- (MER- 0262) – IRED Scraping

  • As of GMT 062, Saturday 3/ 3/ 01, the FORE canister on the IRED was considered hard-
    failed. It is assumed that the side- to- side movement of the flexpacks caused them to
    scrape against the canister housing, caused the flexpack spokes to break rendering
    the entire FORE Canister unusable.

  • The next resupply of the IRED canisters occurs on 6A, so no change out of the
    canisters can occur until that time. Also, there are no spare flexpacks onboard to
    affect repair.

    CDRA Air Save Pump

  • CDRA Air Save Pump not responding during ABIT of the system. Pump and cable to
    be flown on 5A. 1 PRACA PR# 2506 was generated to specifically track this
    nonconformance.

    Node 1 PMA Duct Missing Parts.

  • Duct tape was used. Under investigation. A PRACA is expected.

    SKV 2 Shutdown Due to Oil Pressure Sensor.

  • At approximately GMT 2001/ 065: 09: 00, the Service Module (SM) SKV 2 unit (air
    conditioner) shut down because of low oil pressure.
  • It is thought that the problem is the oil pressure sensor. There is no plan to restart
    CKB2. The plan is to replace the power block on SKV1 and operate it. Short- term
    solution is to install the power supply to SKV1 and activate SKV1 to re gain cooling in
    the SM. There are no further plans yet for long- term solution for fixing SKV2.
    Compressor will probably need to be changed out, but no words from Moscow on this
    yet.

    Unexpected momentum/ attitude disturbances.

  • “At ~GMT47/ 17: 54, momentum was seen to increase at least 3000Nms in roll, in
    5000Nms in yaw, and 1200Nms in pitch.
  • Initial attitude deviations were seen as high as -12 degrees in roll and 2 degrees in
    yaw in pitch. Since GNC was previously in Momentum Manager, we saw oscillations
    in attitude and momentum until we hit the LOS period.
  • At the same time, a PMA2 depressurization through the lab pressure control
    assembly non- propulsive vent was initiated. This may be related to the momentum
    and attitude disturbances.”

    EEATCS Radiator deploy indicator

  • At GMT 2001/ 031: 09: 43 the ‘Deployed’ limit switch transitioned from ‘yes’ to ‘no’. This
    event signaled the start of a 90minute cycle in which the switch indicates ‘no’ for 40-
    50 minutes and then ‘yes’ for 40- 50 minutes repeating continuously. Initially (GMT
    day 031) the ‘yes’/ ‘no’ break was approximately 50/ 40 minutes. Currently (GMT 044)
    the ‘yes’/ ‘no’ break is around 40/ 50 minutes. Recommend that this be converted to a
    PRACA. This is related to PRACA 2437.

    CPA connector off or loose on Node 1 Nadir CPA- 1 (PMA- 3 I/ F).

    J1 secondary power connector on the Node 1 Nadir CBM CPA- 3 was found damaged
    during 5A. The connector appears to have been damaged by crew contact during
    transfer operations from Orbiter into ISS through PMA- 3. CPA R& R was performed
    on 3/ 1/ 01, followed by successful checkout of the CBM. The damaged CPA was
    stowed with the remaining 3 CPAs from the Node Forward CBM, and is scheduled for
    return on 6A. Chit ISS0136 was written by MER Structures & Mechanisms to keep one
    of the Node 1 fwd CPAs on board ISS in order to replace the broken CPA.
    Subsequently Chit ISS0147 was written, asking for all four Node 1 fwd CPA to be left
    on board as spares. All four Node 1 fwd CPAs were left on board and the Node 1
    Nadir CPA3 was removed and replaced with the Node 1 Fwd CPA3 by the increment
    crew. The CBM was subsequently powered and checked out nominally.

    Seat Track Fit Problem

    OSO reported that the crew experienced problems fitting studs in the LAB Seat
    Tracks. The studs in question do not seem to fit any of the LAB Seat Tracks. Item
    involved is a GFE stud (SDG- 33110621- 001 – a utility stud with a tie- ring on one end)
    that is not necessarily compatible with Boeing seat tracks.

    BGA 4B Observed vs Commanded State Failure

    During Progress 2P undocking on GMT 039, a series of three 4B BGA Observed vrs
    Commanded State Failure were observed. The first and third C& Ws were caused by
    Progress thruster firing and the second was caused by the ISS Vehicle regaining
    attitude control after Progress undocking.

    ISSMER
    Ammonia QDs

    1) M3 QD Leak on Z1 Tray

    2) F1 and F3 detemp buttons hard to depress on Z1 tray

    3) Z1 tray QD Caps stuck and could not be removed (M3, M7)

    4) QD caps on Z1 Active Male QDs did not lock (M1,M3, M7)

    Photos taken on EVA 3 and are under review.

    EV& CS input. Since the caps that were installed on M3, M5 and M7 were easier to
    remove on EVA 3 and two of the three were installed to the lock position, we believe
    that the difficulty in removing the caps from Z1 could have been the result of a bad
    work site during EVA 1. During EVA3, PMA2 was not installed on Z1 and the work
    site improved. The difficulty could have also have been due to the fact that the caps
    are spring loaded and it does take some amount of force to install them. A specific
    tool was built for install and remove the caps to assist in increasing the torque. The
    5A crew did not use the tool for this operation.

    Aluminized Kapton Bubbles on 7th radiator panel.

    Bubbles observed in the alum. kapton tape on the end of the 7th panel of stowed
    radiator. Requested via CHIT EVA photos be taken to ensure we are good to deploy
    radiator. Need to review available video/ photo. An initial assessment shows that the
    bubbling seen on one side of the STCR can degrade the freeze protection
    performance but there is still positive margin. Detailed thermal analysis is currently
    underway to evaluate worst- case scenarios. LMMFC will perform coupon tests to try
    to recreate bubbling. A PRACA is expected.

    External TCS Fluid caps out of config.

    Caps installed on two of four Z1 jumper male QDs on the Z1 tray after Z1 to Lab QD
    connections were made during EVA1 were reported to not be screwed all the way
    down. On EVA3, one was corrected, but a second (F7?) was still reported to not be
    screwed all the way down. Z1 jumper male QDs are intended to be used a gain in the
    future when the EEATCS is disconnected from the Lab on 12A. 1 and the EATCS is
    activated. The Z1 to Lab females will be re- installed on the Z1 jumpers to permit
    circulation of NH3 in the decommissioned EEATCS. No corrective action is required
    for the Z1 jumper male QD caps. Continued fit checks will screen out ground
    damaged or contaminated caps or male QDs (causes 2 & 3).

    EEATCS Radiator Return Temperature Difference.

    During Stage 4A EEATC operations differences have been observed between the
    Loop A and Loop B radiator return temperatures. Outlet temps not consistent with
    analysis (outlet temperatures greater than predictions).

    SSC Network not Functioning

  • The crew has not been able to set up the SSC network. Printer problem turned out to
    be a cabling problem. Crew resolved problem. Reference OCA message 335.
    Network became operational after replacing a network card and the OCA laptop on 4A.
    The IFI will be closed once a FIAR is generated. Some assistance may be required
    from the ISS crew to retrieve the serial number of the PCMCIA card (bagged and
    tagged by Shep) for FIAR documentation.
  • Printer started working after network cable was reconfigured by the crew.

  • SpaceRef staff editor.